# ASKAP and VLASS search for a radio-continuum counterpart of ultra-high-energy neutrino event KM3–230213A

### v1

@@ Abstract
We present the results of an Australian Square Kilometre Array Pathfinder (ASKAP) 944 MHz and Very Large Array Sky Survey (VLASS) 3 GHz search for a radio-continuum counterpart of the recent ultra-high-energy (UHE) neutrino event, KM3–230213A. Using ASKAP, we catalog 1052 radio sources within the 15 radius search area (68% certainty region) around the particle’s calculated origin, 10 of which we classify as blazar candidates based on their radio spectra. The most prominent radio source in the search area is the nearby spiral galaxy UGCA 127 (nicknamed Phaedra111From Greek: , a Cretan princess of Greek Mythology, derived from Phaidros, Greek: , meaning ’bright’.). Its non-thermal radio spectrum classifies it as a non-blazar active galactic nucleus (AGN). We also present an extended radio source, WISEA J061715.89075455.4 (nicknamed Hebe222From Greek: H, the Greek goddess of youth.), located only 7′ from the geometric center of the search area, with a very unusual highly polarized compact component. Finally, we present a strong radio source, EMU J062248072246 (nicknamed Narcissus333From Greek was a self-absorbed hunter from Thespiae in Boeotia.), which has a maximum self-absorption spectral slope of +2.5 at low frequencies, and exhibits flux density variability over the 5-year VLASS 3 GHz survey.
@@ 1 Introduction
Several origins for KM3–230213A have been investigated, including Galactic, which was deemed unlikely (Adriani et al., 2025), a GRB origin (Amelino-Camelia et al., 2025), a cosmogenic origin (The KM3NeT collaboration, 2025; KM3NeT Collaboration et al., 2025), and a blazar origin, specifically the well-known blazar, PMN J0606–0724, which experienced a radio flare within five days of the KM3–230213A event (Collaboration et al., 2025). We here present the results of a search for a potential radio-continuum counterpart of the KM3–230213A event using the ASKAP- Evolutionary Map of the Universe (EMU) survey and the Quick-Look images from VLASS (Lacy et al., 2020). The data used in this paper are explained in Appendix A.
@@ 2.2 Objects of interest
We find several objects of interest. The potential blazar counterpart, PMN J0606–0724, analyzed by Collaboration et al. (2025), is not discussed here as it is located outside of the EMU tile coverage.
@@ 2.2.2 Phaedra (UGCA 127; EMU J062056082949)
We generate a spectral index map with high-resolution EMU, MeerKAT, and 3-epoch averaged VLASS data using the miriad (Sault et al., 1995) maths function. We convolve all images to a common beam size of 10″10″and calculate the spectral index per pixel. The VLASS image is only reliable for the inner emission due to missing short spacing, so the images were cut at a 2 mJy beam-1 level. We measure an average over the entire region shown (Figure 2, middle left), with the west component being flattest . This is consistent with a central engine of higher energy injection and particle acceleration in the west component.
@@
Phaedra is unlikely to be a blazar due to its steeper spectral index and lack of detected variability. However, it hosts several potential areas of UHE particle acceleration. The proposed AGN jets are a potential origin, but they appear oriented away from us, and the star formation rate (SFR) is inconclusive. These jet orientations have been observed in other galaxies, some of which are possible high-energy neutrino sources, e.g. NGC 1068 (Padovani et al., 2024), NGC 7469 (Sommani et al., 2024), and NGC 5938 (Zakir et al. in prep). It is possible that these jets are emitting UHE-CRs into surrounding space; by interacting with ambient gas or photons, they can produce neutrinos. This raises the possibility that the neutrino direction may not point directly to the particle acceleration site.
In order to assess the probability of chance association of Phaedra we used the bright galaxy number counts (Yasuda et al., 2001), which implies 0.4 galaxies with a similar or brighter magnitudes in a 15 radius. This can mean a likely chance association, but we note that galaxies with radio properties similar to that of Phaedra are rarer than the total number counts.
@@
The closest extended radio source to the target position is found 7′ to its south. Called Hebe here, it is a double radio galaxy with diffuse lobes and an extraordinary slightly resolved polarized component, offset by 3″ along the direction of its major axis from the IR galaxy WISEA J061715.89075455.4 (see Figure 3). At a redshift of = 0.1250.017 (Bilicki et al., 2014), this offset corresponds to 8 kpc. No compact component is visible in the total intensity image.
The bright peak polarized intensity is 0.8 mJy/(18″ beam-1) above its local background. To estimate its fractional polarization, we fit three Gaussians to a one-dimensional cut through the total intensity image and forced it to fit one component at the polarized peak position. This yielded a flux density of 1.50.13 mJy/beam, corresponding to a fractional polarization of . Even core-dominated quasars have typical fractional polarizations of only a few percent (Pushkarev et al., 2023), although the resolved components of pc-scale Very Long Baseline Interferometry (VLBI) jets can reach levels of 10s of percent (e.g., Jorstad et al., 2005). At 8 kpc from the presumed AGN core in WISEA J061715.89075455.4, Hebe ’s central polarized component could also be a jet, which is only slightly resolved here, although its direction would be almost perpendicular to the more extended emission.
@@
It is unclear whether Hebe is related to the class of blazars or other core-dominated radio AGN. Blazars are of interest as possible sources of cosmic neutrinos (Albert et al., 2024), so further investigation of Hebe is warranted. Monitoring in radio polarization, and searches for possible optical polarization in WISEA J061715.89075455.4 would be especially valueable.
@@ 2.2.4 VLASS investigation and Narcissus (EMU J062248072246)
The most obviously variable source across the VLASS epochs (2017 November 30, 2020 October 9, and 2023 March 10) is the southern source in Figure 4, (EMU J062248072246, hereinafter Narcissus). At the native VLASS 3″ resolution, its peak flux densities are 85.4, 70.4, and 65.2 mJy beam-1 with errors of 0.15 mJy beam-1. The northern source (EMU J062248072246) also appears variable, with peak flux densities of 2.1, 3.7 and 1.3 mJy beam-1. Both components are consistent with being unresolved within the errors. We also search within the EMU data, but detect no variability for Narcissus during the 5-hr observation. Both sources have WISE counterparts (WISEA J062244.90–072334.9 – south, and WISEA J062248.41–072247.9 – north), with colors consistent with AGNs, as expected for radio-selected samples (e.g., Wong et al., 2025).
We combine 4.85 GHz data (Griffith et al., 1995) with our catalog data to measure the broadband spectrum of Narcissus (Figure 4, left). The spectrum shows a maximum possible low-frequency slope of , indicating a homogeneous source with synchrotron self-absorption. At the high, optically thin part of the spectrum, where the variability is observed, the slope is =–0.5, consistent with the limit for strong (non-relativistic) Fermi I shock acceleration (Filipović & Tothill, 2021).
Narcissus differs from many blazars studied with VLBI, e.g., where varying optical depths due to an expanding jet have long been known to flatten the spectrum (O’Dell, 1988). In the low-frequency, absorbed part of the spectrum, inhomogeneities will also flatten the slope, as is observed (e.g., Collier et al., 2018; Kim et al., 2022). For Narcissus, homogeneity would be required over the long time scales spanned by the observations. Further multi-frequency monitoring and VLBI observations are necessary to separate the spectra of the short- and long-term components.
@@ 3 Conclusion
We catalog ASKAP and VLASS radio sources in the 68% confidence region of the UHE neutrino event, KM3–230213A. We identify ten possible blazar candidates in the field that could produce UHE neutrinos. We further analyze three individual objects, the nearby spiral galaxy UGCA 127 (Phaedra), the radio galaxy (Hebe) closest to the geometric search area center and an unusual compact radio source (Narcissus). Phaedra is the most prominent radio source in the field and is a star-forming galaxy. However, jets may be interacting with the galaxy plane itself, causing starburst regions. This possibility, as well as that of SNRs and microquasars within the galaxy, and the well-known XRB, provide areas of possible neutrino generation within Phaedra. Hebe contains an unusual, very highly polarized compact component, likely associated with one of a triplet of galaxies in a common envelope. Narcissus is a radio source that displays radio variability in VLASS and a spectrum consistent with a blazar or a self-absorbed synchrotron source. These properties make Narcissus a strong candidate for the neutrino origin. All of these systems deserve further investigation as possible origins for KM3–230213A.
@@ 4 Acknowledgments
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara / the Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. The Australian Commonwealth Scientific and Industrial Research Organisation (CSIRO)’s ASKAP radio telescope is part of the Australia Telescope National Facility555https://ror.org/05qajvd42. Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. We thank C.W. James for discussion relating to the origin of the neutrino event and Yuri Y. Kovalev for the radio-continuum discussion. This research was funded in part by the National Science Centre, Poland, grant Nos.2018/30/E/ST9/00208 and 2023/49/B/ST9/00066. This research was supported by the Sydney Informatics Hub (SIH), a core research facility at the University of Sydney. This work was also supported by software support resources awarded under the Astronomy Data and Computing Services (ADACS) Merit Allocation Program. ADACS is funded from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government and managed by Astronomy Australia Limited (AAL).

### v2

@@ Abstract
We present the results of an Australian Square Kilometre Array Pathfinder (ASKAP) 944 MHz and Very Large Array Sky Survey (VLASS) 3 GHz search for a radio-continuum counterpart of the recent ultra-high-energy (UHE) neutrino event, KM3–230213A. Using ASKAP, we catalog 1052 radio sources within the 15 radius search area (68% certainty region) around the particle’s calculated origin, 10 of which we classify as blazar candidates based on their radio spectra. The most prominent radio source in the search area is the nearby spiral galaxy UGCA 127 (nicknamed Phaedra111From Greek: , a Cretan princess of Greek Mythology, derived from Phaidros, Greek: , meaning ’bright’.). Its non-thermal radio spectrum classifies it as a non-blazar active galactic nucleus (AGN). We also present an extended radio source, WISEA J061715.89075455.4 (nicknamed Hebe222From Greek: H, the Greek goddess of youth.), located only 7′ from the geometric center of the search area, with a very unusual highly polarized compact component. Finally, we present a strong radio source, EMU J062248072246 (nicknamed Narcissus333From Greek was a self-absorbed hunter from Thespiae in Boeotia.), which can be modeled either with a synchrotron self-absorbed spectrum or synchrotron emission from thermal electrons. It exhibits flux density variability over the 5-year VLASS 3 GHz survey.
@@ 1 Introduction
Several origins for KM3–230213A have been investigated, including Galactic, which was deemed unlikely (Adriani et al., 2025), a GRB origin (Amelino-Camelia et al., 2025), a cosmogenic origin (The KM3NeT collaboration, 2025; KM3NeT Collaboration et al., 2025), and a blazar origin, specifically the well-known blazar, PMN J0606–0724, which experienced a radio flare within five days of the KM3–230213A event (KM3NeT Collaboration et al., 2025). We here present the results of a search for a potential radio-continuum counterpart of the KM3–230213A event using the ASKAP- Evolutionary Map of the Universe (EMU) survey and the Quick-Look images from VLASS (Lacy et al., 2020). The data used in this paper are explained in Appendix A.
@@ 2.2 Objects of interest
We find several objects of interest. The potential blazar counterpart, PMN J0606–0724, analyzed by KM3NeT Collaboration et al. (2025), is not discussed here as it is located outside of the EMU tile coverage.
@@ 2.2.2 Phaedra (UGCA 127; EMU J062056082949)
We generate a spectral index map with high-resolution EMU, MeerKAT, and 3-epoch averaged VLASS data using the miriad (Sault et al., 1995) maths function. We convolve all images to a common beam size of 10″10″ and calculate the spectral index per pixel. The VLASS image is only reliable for the inner emission due to missing short spacing, so the images were cut at a 2 mJy beam-1 level. We measure an average over the entire region shown (Figure 2, middle left), with the west component being the flattest . This is consistent with a central engine of higher energy injection and particle acceleration in the west component.
@@
Phaedra is unlikely to contain a blazar due to its steeper spectral index and lack of detected variability. However, it hosts several potential areas of UHE particle acceleration. Due to particle escape as in the commonly used “leaky box” models, the spectrum of the diffuse CRs within spiral galaxies is likely different than the CR spectra injected from the accelerators. Furthermore, reacceleration of CRs, such as for example in the so-called “espresso” model, where the CRs produced at other locations of the galaxy are reaccelerated within the jets of a central AGN engine, can also change the resulting CR distribution and contribute to the production of UHECRs (Caprioli, 2015). These processes might be identified by differentiating areas of different spectral indices within a galaxy, such as done by Duric et al. (1995); Duric (2000). However, our radio data is unfortunately not of sufficiently high resolution for such a detailed spectral analysis. The proposed AGN jets are a potential origin, but they appear oriented away from us, and the star formation rate (SFR) is inconclusive. These jet orientations have been observed in other galaxies, some of which are possible high-energy neutrino sources, e.g. NGC 1068 (Padovani et al., 2024), NGC 7469 (Sommani et al., 2024), and NGC 5938 (Zakir et al. in prep). It is possible that these jets are emitting UHE-CRs into surrounding space; by interacting with ambient gas or photons, they can produce neutrinos. This raises the possibility that the neutrino direction may not point directly to the particle acceleration site.
To assess the probability of chance association of Phaedra, we used the bright galaxy number counts (Yasuda et al., 2001), which implies 0.4 galaxies with a similar or brighter magnitude in a 15 radius. This can mean a likely chance association, but we note that galaxies with radio properties similar to that of Phaedra are rarer than the total number counts.
@@
The closest extended radio source to the target position is found 7′ to its south. Called Hebe here, it is a double radio galaxy with diffuse lobes and an extraordinary slightly resolved polarized component, offset by 3″ along the direction of its major axis from the IR galaxy WISEA J061715.89075455.4 (see Figure 3). At a redshift of = 0.1250.017 ( = 553.8938.77 pc; Bilicki et al., 2014), this offset corresponds to 8 kpc. No compact component is visible in the total intensity image, but we made a very rough estimate of its contribution by forcing a fit to a Gaussian component at that position, along with additional components for the lobes. This yielded a flux of 1.5 mJy/(18″ beam).
The bright peak polarized intensity is 0.8 mJy beam-1 above its local background. There is clear depolarization, consistent, e.g., with a Burn-slab model with an rms scatter of . After correction for depolarization, this corresponds to an apparent remarkable 100% polarization, varying only slightly with the unknown spectral index. Even core-dominated quasars have typical fractional polarizations of only a few percent (Pushkarev et al., 2023), although the resolved components of pc-scale Very Long Baseline Interferometry (VLBI) jets can reach levels of 10s of percent (e.g., Jorstad et al., 2005). The rotation measure of the polarized peak is , close to the value of the local Galactic foreground, (; Hutschenreuter et al., 2022), so there is no evidence for a large, additional Faraday component local to Hebe.
At 8 kpc from the presumed AGN core in WISEA J061715.89075455.4, Hebe’s central polarized component could also be a jet, which is only slightly resolved here, although its direction would be almost perpendicular to the more extended emission. We estimate an integrated spectral index for Hebe (see Table 1), although we are not able to measure any spatial-spectral structure, especially of the polarized core.
@@
It is unclear whether Hebe is related to the class of blazars or other core-dominated radio AGN. Blazars are of interest as possible sources of cosmic neutrinos (Albert et al., 2024), so further investigation of Hebe is warranted. One interesting model to explore for Hebe would be the “espresso” scheme (Caprioli, 2015) using relativistic jets for the generation of UHECRs. Monitoring in radio polarization and searching for possible optical polarization in WISEA J061715.89075455.4 would be especially valuable.
@@ 2.2.4 VLASS investigation and Narcissus (EMU J062248072246)
The most obviously variable source across the VLASS epochs (2017 November 30, 2020 October 9, and 2023 March 10) is the southern source in Figure 4, (EMU J062248072246, hereinafter Narcissus). At the native VLASS 3″ resolution, its peak flux densities are 85.4, 70.4, and 65.2 mJy beam-1 with errors of 0.15 mJy beam-1. The northern source (EMU J062248072246) also appears variable, with peak flux densities of 2.1, 3.7 and 1.3 mJy beam-1. Both components are consistent with being unresolved within the errors. We also search within the EMU data, but detect no variability for Narcissus during the 5-hr observation. Both sources have WISE counterparts (WISEA J062244.90–072334.9 – south, and WISEA J062248.41–072247.9 – north), with colors consistent with AGNs, as expected for radio-selected samples (e.g., Wong et al., 2025). To better separate the adjacent bright WISE source in the south, we looked at the unWISE colors, which yield W1–W2=1.3, consistent with an AGN. Approximately 4% of compact sources at 3 GHz exhibit variability at the 30% level (Mooley et al., 2016), and are expected to be associated with AGN, so these results themselves are not unusual.
What is unusual is the broadband spectra of Narcissus, when we combine 4.85 GHz data (Griffith et al., 1995) with our catalog data (Figure 4, left). We show two possible models. The first model, with the maximum possible low-frequency slope of , corresponds to a compact, homogeneous source with synchrotron self-absorption. At the high, optically thin part of the spectrum, where the variability is observed, the slope is =–0.5, consistent with the limit for strong (non-relativistic) Fermi I shock acceleration (Filipović & Tothill, 2021). Although an AGN connection for Narcissus is attractive, we cannot rule out a galactic origin. We therefore show an alternative model in Figure 4, with a low-frequency slope of . This is similar to shock models explored by Margalit & Quataert (2024), e.g., their Fig. 3(b), used to describe explosive events such as SN1998bw and AT2018cow. Thermal absorption is another possibility for the extremely low-frequency slope.
Independent of the origins of Narcissus, Galactic or extragalactic, the steep low-frequency slope indicates that the emission is dominated by a compact, homogeneous component. This is built-in to explosive models relying on a single shock, whereas inhomogeneities, or multiple components, would broaden the spectrum and flatten the slope. However, it is not clear how the need for a compact, homogeneous source applies to Narcissussince the broadband spectra span many years of observations. Further multi-frequency monitoring and VLBI observations are necessary to separate the spectra of the short- and long-term components and to compare, e.g., to the evolutionary scenarios, e.g., of Margalit & Quataert (2021).
@@ 3 Conclusion
We catalog ASKAP and VLASS radio sources in the 68% confidence region of the UHE neutrino event, KM3–230213A. We identify ten possible blazar candidates in the field that could produce UHE neutrinos. We further analyze three individual objects, the nearby spiral galaxy UGCA 127 (Phaedra), the radio galaxy (Hebe) closest to the geometric search area center and an unusual compact radio source (Narcissus). Phaedra is the most prominent radio source in the field and is a star-forming galaxy. However, jets may be interacting with the galaxy plane itself, causing starburst regions. This possibility, as well as that of SNRs and microquasars within the galaxy, and the well-known XRB, provide areas of possible neutrino generation within Phaedra. Hebe contains an unusual, very highly polarized compact component, likely associated with one of a triplet of galaxies in a common envelope. Narcissus is a radio source that displays radio variability in VLASS, originating in a compact, heavily absorbed AGN or a more local explosive event. These properties are consistent with known neutrino produces, making Narcissus a plausible candidate for the neutrino origin. Both Hebe and Narcissus are clearly associated with an explosive activity that could produce UHE neutrinos, making them the most likely candidates. Phaedra shows evidence of more high-energy emission than other galaxies, making it a possible candidate, albeit this evidence is somewhat weaker than for the others. All of these systems deserve further investigation as possible origins for KM3–230213A.
@@ 4 Acknowledgments
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara / the Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. The Australian Commonwealth Scientific and Industrial Research Organisation (CSIRO)’s ASKAP radio telescope is part of the Australia Telescope National Facility555https://ror.org/05qajvd42. Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. We thank C.W. James for discussion relating to the origin of the neutrino event and Yuri Y. Kovalev for the radio-continuum discussion. This research was funded in part by the National Science Centre, Poland, grant Nos.2018/30/E/ST9/00208 and 2023/49/B/ST9/00066. This research was supported by the Sydney Informatics Hub (SIH), a core research facility at the University of Sydney. This work was also supported by software support resources awarded under the Astronomy Data and Computing Services (ADACS) Merit Allocation Program. ADACS is funded from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government and managed by Astronomy Australia Limited (AAL). We thank an anonymous referee for excellent comments and suggestions, which greatly improved this paper.